OpenStudy (anonymous):

10x - 10 -7x 3x - 2 10x - 10 -7x 3x - 2 @Mathematics

hero (hero):

Problem not posted clearly

OpenStudy (anonymous):

\[\ge\] this is what was missing

hero (hero):

Chrissyj, I'm on skype btw. You know how slow skype is

OpenStudy (anonymous):

this symbol \[\ge\] is after the 7x

hero (hero):

\[10x - 10 - 7x \ge 3x - 2\] \[3x - 10 \ge 3x - 2\] -10 is not greater than or equal to -2 I hate equation editor

hero (hero):

If you notice, in that second step, the 3x's cancel leaving only -10 and -2